OpenStudy (anonymous):

5/12m + 2/3 = -3/4 solve for m

OpenStudy (anonymous):

Rearrange the terms so that you can combine it, constants with constants

OpenStudy (anonymous):

what do you mean that is very confusing

OpenStudy (anonymous):

5m / 12 + 2 / 3 = -3 / 4 5m / 12 = -3 / 4 - 2 / 3 = -17 / 12 5m = -17 m = -17 / 5

OpenStudy (anonymous):

Or if the m is below 5 / (12m) + 2 / 3 = -3 / 4 5 / (12m) = -17 / 12 12m = 5 / (-17 / 12) = -60 / 17 m = -60 / 17 / 12 = -5 / 17

OpenStudy (anonymous):

the choices are these though: M=-1/5 M=-5/144 M=-3 2/5 M=-85/144

OpenStudy (anonymous):

-17 / 5 = -3 2 / 5